OpenStudy (anonymous):

Find the period of the function. y = -5 cos x

OpenStudy (anonymous):

@abb0t can you helppp

OpenStudy (anonymous):

I can.

OpenStudy (anonymous):

yay!!!!!!!!!!!!!!!:D

OpenStudy (anonymous):

You have to find \(p\) such that: \[ -5\cos(x) = -5\cos(x+np), \quad n\in \mathbb{N} \]But that is really troublesome.

OpenStudy (anonymous):

We know that \(\cos\)'s period is just the angle of a full revolution.

OpenStudy (anonymous):

So in degree mode it is \(360^\circ\), and in radian mode \(2\pi\), and in revolutions it is \(1\).

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ahh so the answer is 2pi right?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thankyouu